### v3

@@ 5.2. Oriented exchange graphs are Hasse quivers
``$(a)\Rightarrow(b)"$: Suppose that $[{\bf x}_{w},\widetilde{B}_{w}]\rightarrow[{\bf x}_{t},\widetilde{B}_{t}]$ is an arrow in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. Then we know that $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$.
Since the two seeds $({\bf x}_{t},\widetilde{B}_{t})$ and $({\bf x}_{w},\widetilde{B}_{w})$ differ by once mutation, for simplicity, we can assume $({\bf x}_{t},\widetilde{B}_{t})=\mu_{1}({\bf x}_{w},\widetilde{B}_{w})$. If we write ${\bf x}_{w}=(z_{1},z_{2},\ldots,z_{m})$, then ${\bf x}_{t}$ has the form ${\bf x}_{t}=(z_{1}^{\prime},z_{2},\ldots,z_{m})$.
If there exists a seed $({\bf x}_{s},\widetilde{B}_{s})$ such that $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$. By Lemma 5.11, we have $\{z_{2},\ldots,z_{m}\}=[{\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$. Since $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]$, we know that $[{\bf x}_{s}]\neq[{\bf x}_{w}]$. Thus the two clusters ${\bf x}_{s}$ and ${\bf x}_{w}$ have exactly $m-1$ common cluster variables $z_{2},\ldots,z_{m}$. Then by [cao-li-2020, Corollary 3], we get $[{\bf x}_{s},\widetilde{B}_{s}]=[\mu_{1}({\bf x}_{w},\widetilde{B}_{w})]=[{\bf x}_{t},\widetilde{B}_{t}]$. So there exists no seed $({\bf x}_{s},\widetilde{B}_{s})$ such that $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$.
``$(b)\Rightarrow(a)"$: Since $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$, there exists a non-empty sequence $\overleftarrow{\mu}$ of green mutations from $[{\bf x}_{w},\widetilde{B}_{w}]$ to $[{\bf x}_{t},\widetilde{B}_{t}]$. Since there exists no seed $({\bf x}_{s},\widetilde{B}_{s})$ such that
$[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}],$
the length of $\overleftarrow{\mu}$ must be $1$, i.e., $[{\bf x}_{t},\widetilde{B}_{t}]$ is obtained from $[{\bf x}_{w},\widetilde{B}_{w}]$ by once green mutation. So there exists an arrow $[({\bf x}_{w},\widetilde{B}_{w})]\rightarrow[({\bf x}_{t},\widetilde{B}_{t})]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. ∎
@@
The oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$ coincides with the Hasse quiver $\overrightarrow{\mathcal{H}}^{\prime}$ of the poset $(\mathcal{H}_{0},\leq_{t_{0}})$.
@@ 5.3. Dominant sets and torsion classes
In this subsection we briefly explain why the dominant sets can be viewed as an replacement of torsion classes in $\tau$-tilting theory. More discussions will be given in [cao_2025].
We fix a finite dimensional basic algebra $A$ over a field $K$. Denote by $\operatorname{\mathsf{mod}}\nolimits A$ the category of finitely generated right $A$-modules, and by $\tau$ the Auslander-Reiten translation in $\operatorname{\mathsf{mod}}\nolimits A$. The isomorphism classes of indecomposable projective modules in $\operatorname{\mathsf{mod}}\nolimits A$ are denoted by $P_{1},\ldots,P_{n}$.
@@
  • $\prescript{\bot}{}{M}:=\{X\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(X,M)=0\}$.
  • $M^{\bot}:=\{Y\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(M,Y)=0\}$.
@@
Theorem 5.15.
[air_2014, Proposition 1.2 (b) and Theorem 2.7] The following statements hold.
@@
$\text{Cluster algebras}&\hskip 5.69054pt\tau\text{-tilting theory}\\[5.69054pt] \hline\cr&\\[-8.53581pt] \text{Seeds}&\hskip 5.69054pt\tau\text{-tilting pairs}\\[5.69054pt] \hline\cr&\\[-5.69054pt] \text{Mutations of seeds}&\hskip 5.69054pt\text{Mutations of }\tau\text{-tilting pairs}\\[5.69054pt] \hline\cr&\\[-5.69054pt] \text{Initial unfrozen cluster variables }&\hskip 5.69054pt(0,P_{1}),\ldots,(0,P_{n})\\[5.69054pt] \hline\cr&\\[-8.53581pt] \text{Non-initial unfrozen cluster variables}&\hskip 5.69054pt\text{Indecomposable }\tau\text{-rigid modules}\\[5.69054pt] \hline\cr&\\[-8.53581pt] \text{Unfrozen cluster variables}&\hskip 5.69054pt\text{Indecomposable }\tau\text{-rigid pairs}\\[5.69054pt] \hline\cr\end{array}$
@@
$\delta_{M}:=(a_{1}-b_{1},\ldots,a_{n}-b_{n})^{T}\in\mathbb{Z}^{n}$
is called the $\delta$-vector of $M$ and the vector ${\bf g}_{M}:=-\delta_{M}$ is called the $g$-vector of $M$.
@@
$\delta_{(M,P)}:=\delta_{M}-\delta_{P},\;\;{\bf g}_{(M,P)}:=-\delta_{(M,P)}.$
With this definition, one can see that the $g$-vector ${\bf g}_{(0,P_{k})}$ of $(0,P_{k})$ is the $k$th column of $I_{n}$.
@@
where $\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)$ is the quotient module Grassmannian with dimensional vector ${\bf v}$ and $\chi$ is the Euler–Poincaré characteristic.
Since the zero module is a quotient module of $M$, $F_{M}$ is a polynomial with constant term $1$. Thus
$F_{M}[{\bf r}]=\max\{{\bf v}^{T}{\bf r}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M))\neq 0\}\geq 0$
for any ${\bf r}\in\mathbb{Z}^{n}$.
@@
Notice that when we write $U\in\mathcal{X}$, we mean that $U$ is an indecomposable $\tau$-rigid pair, that is, either $U\in\mathcal{X}_{+}$ or $U=(0,P_{k})$ for some $k\in[1,n]$.
Similarly to the dominant sets in cluster algebras, we can define dominant sets in $\tau$-tilting theory.
@@
$\operatorname{\mathsf{dom}}[M,P]$ ::=:= $\{U\in\mathcal{X}\mid F_{U}[{\bf g}_{(M,P)}]=0\},$
$\operatorname{\mathsf{dom}}_{+}[M,P]$ ::=:= $\{U\in\mathcal{X}_{+}\mid F_{U}[{\bf g}_{(M,P)}]=0\}.$
Since $F_{(0,P_{k})}=1$ for $k=1,\ldots,n$, we see that the set $\{(0,P_{1}),\ldots,(0,P_{n})\}$ is always contained in the dominant set $\operatorname{\mathsf{dom}}[M,P]$ for a $\tau$-tilting pair. Thus we have
@@
For a module $X\in\operatorname{\mathsf{mod}}\nolimits A$ and a vector $\delta\in\mathbb{R}^{n}$, we write $\delta(X)\in\mathbb{R}$ for the inner product of $\delta$ and the dimensional vector of $X$.
Lemma 5.19.
Denote by
$\overline{\mathcal{T}}_{\delta}=\{N\in\operatorname{\mathsf{mod}}\nolimits A\mid\text{for any quotient module X of }N,\;\;\delta(X)\geq 0\}$
for $\delta\in\mathbb{R}^{n}$. The following statements hold.
@@
    [BKT-2014]*Proposition3.1 The subcategory $\overline{\mathcal{T}}_{\delta}$ is a torsion class in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
    [asai-2021, Proposition 3.11] Let $(M,P)$ be a basic $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$ and $\delta=\delta_{(M,P)}$ the $\delta$-vector of $(M,P)$. Then $\operatorname{\mathsf{Fac}}M=\overline{\mathcal{T}}_{\delta}$.
@@
Lemma 5.20.
[fei_2019a, Theorem 1.4] Let $U\in\operatorname{\mathsf{mod}}\nolimits A$ and let
@@
be the $F$-polynomial of $U$. Then the convex hull $\mathcal{N}(F_{U})$ of $\{{\bf v}\in\mathbb{N}^{n}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U))\neq 0\}$ and the convex hull $\mathcal{N}(U)$ of the dimensional vectors of the quotient modules of $U$ are the same.
@@
(i) Let $\delta=\delta_{(M,P)}$ be the $\delta$-vector of $(M,P)$ and ${\bf g}={\bf g}_{(M,P)}=-\delta$ the $g$-vector of $(M,P)$. Let $U$ be an indecomposable $\tau$-rigid module and
@@
$\max\{{\bf g}(X)\mid X\text{ is a quotient module of }U\}=0,$
equivalently, ${\bf g}(X)\leq 0$ for any quotient module $X$ of $U$. This is also equivalent to $\delta(X)=-{\bf g}(X)\geq 0$ for any quotient module $X$ of $U$. Then by Lemma 5.19, this is equivalent to $U\in\operatorname{\mathsf{Fac}}M$. Hence, we have
@@
Since $\operatorname{\mathsf{Fac}}M$ is the smallest torsion classes containing all the indecomposable $\tau$-rigid modules in $\operatorname{\mathsf{Fac}}M$, we know that $\operatorname{\mathsf{Fac}}M$ is uniquely determined by the dominant set $\operatorname{\mathsf{dom}}[M,P]$. Thus we can view the dominant sets in cluster algebras as a kind of replacement of torsion classes in $\tau$-tilting theory.
@@
The oriented exchange graph of $\tau$-tilting pairs are given as follows:

### v4

@@ 7.2. Oriented exchange graphs are Hasse quivers
$(a)\Rightarrow(b)$": Suppose that ${\bf x}_{w},B_{w}]\rightarrow[{\bf x}_{t},B_{t}]$ is an arrow in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. Then we know that ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{t},B_{t}]$.
Since the two seeds $({\bf x}_{t},B_{t})$ and $({\bf x}_{w},B_{w})$ differ by once mutation, for simplicity, we can assume $({\bf x}_{t},B_{t})=\mu_{1}({\bf x}_{w},B_{w})$. If we write ${\bf x}_{w}=(z_{1},z_{2},\ldots,z_{n})$, then ${\bf x}_{t}$ has the form ${\bf x}_{t}=(z_{1}^{\prime},z_{2},\ldots,z_{n})$.
If there exists a seed $({\bf x}_{s},B_{s})$ such that ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]\leq_{t_{0}}[{\bf x}_{t},B_{t}]$. Then we have $\operatorname{\mathsf{dom}}^{t_{0}}[w]\subsetneq\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$. By Lemma 7.12, we have $\{z_{2},\ldots,z_{m}\}=[{\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$. Since ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]$, we know that ${\bf x}_{s}]\neq[{\bf x}_{w}]$. Thus the two clusters ${\bf x}_{s}$ and ${\bf x}_{w}$ have exactly n1n-1 common cluster variables $z_{2},\ldots,z_{n}$. Then by [cao-li-2020, Corollary 3], we get ${\bf x}_{s},B_{s}]=[\mu_{1}({\bf x}_{w},B_{w})]=[{\bf x}_{t},B_{t}]$. So there exists no seed $({\bf x}_{s},B_{s})$ such that ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]<_{t_{0}}[{\bf x}_{t},B_{t}]$.
$(b)\Rightarrow(a)$": Since ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{t},B_{t}]$, there exists a non-empty sequence $\overleftarrow{\mu}$ of green mutations from ${\bf x}_{w},B_{w}]$ to ${\bf x}_{t},B_{t}]$. Since there exists no seed $({\bf x}_{s},B_{s})$ such that
${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]<_{t_{0}}[{\bf x}_{t},B_{t}],$
the length of $\overleftarrow{\mu}$ must be $1$, i.e., ${\bf x}_{t},B_{t}]$ is obtained from ${\bf x}_{w},B_{w}]$ by once green mutation. So there exists an arrow $({\bf x}_{w},B_{w})]\rightarrow[({\bf x}_{t},B_{t})]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. ∎
@@
[…]
@@ 7.3. Dominant sets and torsion classes
In this subsection we briefly explain why the dominant sets can be viewed as a replacement of torsion classes in $\tau$-tilting theory.
We fix a finite dimensional basic algebra $A$ over a field $\mathbf{k}$. Denote by $\operatorname{\mathsf{mod}}\nolimits A$ the category of finitely generated left $A$-modules, and by $\tau$ the Auslander-Reiten translation in $\operatorname{\mathsf{mod}}\nolimits A$. The isomorphism classes of indecomposable projective modules in $\operatorname{\mathsf{mod}}\nolimits A$ are denoted by $P_{1},\ldots,P_{n}$.
@@
  • $\prescript{\bot}{}{M}\coloneqq\{X\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(X,M)=0\}$.
  • $M^{\bot}\coloneqq\{Y\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(M,Y)=0\}$.
@@
Theorem 7.16 ([air_2014, Prop. 1.2 (b), Thm. 2.7]).
The following statements hold.
@@
[…]
@@
$\delta_{M}\coloneqq(a_{1}-b_{1},\ldots,a_{n}-b_{n})^{T}\in\mathbb{Z}^{n}$
is called the $\delta$-vector of $M$ and the vector ${\bf g}_{M}\coloneqq-\delta_{M}$ is called the $g$-vector of $M$.
@@
$\delta_{(M,P)}\coloneqq\delta_{M}-\delta_{P},\;\;{\bf g}_{(M,P)}\coloneqq-\delta_{(M,P)}.$
With this definition, we can see that the $g$-vector ${\bf g}_{(0,P_{k})}$ of $(0,P_{k})$ is the $k$th column of $I_{n}$.
@@
where $\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)$ is the quotient module Grassmannian of $M$ with dimension vector $\bf v$} and $\chi$ is the Euler–Poincaré characteristic.
Since the zero module is a quotient module of $M$, the polynomial $F_{M}$ has constant term $1$. So we have
$F_{M}[{\bf r}]=\max\{{\bf v}^{T}{\bf r}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M))\neq 0\}\geq 0,\quad\forall{\bf r}\in\mathbb{Z}^{n}.$
@@
Remark 7.19.
Notice that in Section 6 we define $F$-polynomial using submodules (see Definition 6.2), while in this section we define $F$-polynomial using quotient modules. The reason is that in this section we want to work with torsion classes (corresponding to $\tau$-tilting theory) rather than torsion-free classes (corresponding to $\tau^{-1}$-tilting theory). So the algebras/quivers in Section 6 correspond to the opposite algebras/quivers in this section. For example, in Section 6 we use the representations of the quiver $1\leftarrow 2$ to model the cluster algebra $\mathcal{A}$ with initial mutation matrix $B=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$, while in this section we would use the representations of the quiver $1\rightarrow 2$ to model $\mathcal{A}$.
@@
When we write $U\in\mathcal{X}$, we mean that $U$ is an indecomposable $\tau$-rigid pair, that is, either $U\in\mathcal{X}_{+}$ or $U=(0,P_{k})$ for some $k\in[1,n]$.
Following the definition of the dominant sets in cluster algebras, we define the dominant sets in $\tau$-tilting theory.
@@
$\operatorname{\mathsf{dom}}[M,P]$ $\coloneqq\{U\in\mathcal{X}\mid F_{U}[{\bf g}_{(M,P)}]=0\},$
$\operatorname{\mathsf{dom}}_{+}[M,P]$ $\coloneqq\{U\in\mathcal{X}_{+}\mid F_{U}[{\bf g}_{(M,P)}]=0\}.$
Since $F_{(0,P_{k})}=1$ for $k=1,\ldots,n$, the set $\{(0,P_{1}),\ldots,(0,P_{n})\}$ is contained in the dominant set $\operatorname{\mathsf{dom}}[M,P]$ for any $\tau$-tilting pair $(M,P)$. Thus we have
@@
For a module $X\in\operatorname{\mathsf{mod}}\nolimits A$ and a vector $\delta\in\mathbb{R}^{n}$, we write $\delta(X)=\langle\delta,\underline{\dim}X\rangle\in\mathbb{R}$ for the inner product of $\delta$ and the dimension vector of $X$.
Lemma 7.21 ([BKT-2014]*Prop. 3.1, [asai-2021]*Prop. 3.11).
Let $\delta\in\mathbb{R}^{n}$ and set
$\overline{\mathcal{T}}_{\delta}:=\{N\in\operatorname{\mathsf{mod}}\nolimits A\mid\text{for any quotient module X of }N,\;\;\delta(X)\geq 0\}.$
The following statements hold.
@@
    The subcategory $\overline{\mathcal{T}}_{\delta}$ is a torsion class in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
    Let $(M,P)$ be a basic $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$ and $\delta=\delta_{(M,P)}$ the $\delta$-vector of $(M,P)$. Then $\operatorname{\mathsf{Fac}}M=\overline{\mathcal{T}}_{\delta}$.
@@
Lemma 7.22 ([fei_2019a, Theorem 1.4]).
Let $U\in\operatorname{\mathsf{mod}}\nolimits A$ and let
@@
be the $F$-polynomial of $U$. Then the convex hull $\mathcal{N}(F_{U})$ of $\{{\bf v}\in\mathbb{N}^{n}\mid\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(U))\neq 0\}$ and the convex hull $\mathcal{N}(U)$ of the dimension vectors of the quotient modules of $U$ are the same.
@@
(i) Let $\delta=\delta_{(M,P)}$ be the $\delta$-vector of $(M,P)$ and ${\bf g}={\bf g}_{(M,P)}=-\delta$ the $g$-vector of $(M,P)$. Let $U$ be an indecomposable $\tau$-rigid module in $\operatorname{\mathsf{mod}}\nolimits A$ and
@@
$\max\{\langle{\bf g},\underline{\dim}X\rangle\mid X\text{ is a quotient module of }U\}=0,$
equivalently, $\langle{\bf g},\underline{\dim}X\rangle\leq 0$ for any quotient module $X$ of $U$. This is also equivalent to $\delta(X)=-\langle{\bf g},\underline{\dim}X\rangle\geq 0$ for any quotient module $X$ of $U$. Then by Lemma 7.21, this is equivalent to $U\in\operatorname{\mathsf{Fac}}M$. Hence, we have
$\operatorname{\mathsf{dom}}_{+}[M,P]=\mathcal{X}_{+}\cap\operatorname{\mathsf{Fac}}M=\{\text{indecomposable $\tau$-rigid modules in $\operatorname{\mathsf{Fac}}M$}\}.$
@@
Since $\operatorname{\mathsf{Fac}}M$ is the smallest torsion class containing all the indecomposable $\tau$-rigid modules in $\operatorname{\mathsf{Fac}}M$, we know that $\operatorname{\mathsf{Fac}}M$ is uniquely determined by the dominant set $\operatorname{\mathsf{dom}}[M,P]$. Thus we can view the dominant sets as a kind of replacement of torsion classes in $\tau$-tilting theory.
@@
The oriented exchange graph of $\tau$-tilting pairs is as follows: